OpenStudy (anonymous):

Find an exact solution for route5 - 1/ x = route5/2 Then find the approximate solution to the nearest tenth

OpenStudy (primeralph):

could you write it clearly?

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Hello?

OpenStudy (primeralph):

yeah

OpenStudy (anonymous):

Do you have to cross multiply?

OpenStudy (primeralph):

exactly

OpenStudy (anonymous):

So you would get \[2\sqrt{5}-2/\sqrt{5}=\sqrt{5}x\]

OpenStudy (anonymous):

How would you do the rest.

OpenStudy (primeralph):

divide all by sqrt(5)